Two cyclists begin traveling in the same direction on the same bike path. One travels at 20 miles per hour, and the other travels at 14 miles per hour. After how much time will the cyclists be 15 miles apart?
I have NO clue or idea what to do..I don't know any steps or anything. Please help me..thank you.

Answer by CharlesG2(834)

Two cyclists begin traveling in the same direction on the same bike path. One travels at 20 miles per hour, and the other travels at 14 miles per hour. After how much time will the cyclists be 15 miles apart?
I have NO clue or idea what to do..I don't know any steps or anything. Please help me..thank you.
D distance = R rate * T time
Cyclist A goes 20 mph
Cyclist B goes 14 mpg
Distance A travels - Distance B travels = 15 miles
20 * T - 14 * T = 15
6 * T = 15
T = 15/6 = 5/2 = 2.5 hours
Cyclist A goes 20mph * 2.5 hours = 50 miles
Cyclist B goes 14mph * 2.5 hours = 35 miles
50 miles - 35 miles is 15 miles apart in 2.5 hours